WITH THE LATEST VERSION OF PYTHON:

A babysitter charges $2.50 an hour until 9:00 PM when the rate drops to $1.75 an hour (the children are in bed). Write a program that accepts a starting time and ending time in hours and minutes and calculates the total babysitting bill. You may assume that the starting and ending times are in a single 24-hour period. Partial hours should be appropriately prorated.
Answer #1

The code in image:

Code:

print("enter starting time hours")   
st_hours=int(input()) #starting time hours
print("minutes")
st_minutes=int(input()) #starting time minutes

print("enter ending time hours")
et_hours=int(input()) #ending time hours
print("minutes")
et_minutes=int(input()) #ending time minutes

cost=2.50

if (st_hours<21 and et_hours>21): #if starting time is less than 21 and ending time is greater than 21
   a_hours=et_hours-21
   a_minutes=et_minutes #taking time after 21 from ending time ending time-21

   time_minutes=60-st_minutes #converting all the time into minutes
   time_hours=21-(st_hours+1)
   time_hours=time_hours*60
   time=time_hours+time_minutes

   cost=(cost/60)*time
   time=(a_hours*60)+a_minutes
   cost=cost+(1.75/60)*time

   print("cost=$",cost)

elif(st_hours>=21): #if starting time is greater than 21
   time_hours=et_hours-(st_hours+1)
   time_minutes=(60-st_minutes)+et_minutes #converting time into minutes
   time=(time_hours*60)+time_minutes
   cost=(1.75/60)*(time)
   print("cost=$",cost)

elif(et_hours<=21): #if ending time is less than 21
   time_hours=et_hours-(st_hours+1)
   time_minutes=(60-st_minutes)+et_minutes
   time=(time_hours*60)+time_minutes
   cost=(2.50/60)*time
   print("cost=$",cost)

  
  

  
Output:

Here I divided the program into three conditions

(1) If starting time is less than 21 and ending time is greater than 21

Then I converted the time into minutes before 21

ex:- 8:30

21-9= 12*60= 720 + 30 = 750

Then I calculated the bill.

750 * (2.50 / 60)

Then I converted the time after 21 and then calculated the bill and printed it.

(2)If starting time is greater than 21

Converted time into minutes and multiplied it with (1.75 / 60) to get the bill.

(3) If ending time is less than 21

Converted time into minutes and multiplied it with (2.50 / 60) to get the bill.